\section*{References}
\begin{thebibliography}{9}

%\cite{Heisenberg:1935qt}
\bibitem{Heisenberg:1935qt} 
  W.~Heisenberg and H.~Euler,
  %``Consequences of Dirac's theory of the positron,''
  Z.\ Phys.\  {\bf 98}, 714 (1936),
  an English translation is available at [physics/0605038].
  %%CITATION = PHYSICS/0605038;%%
  %%CITATION = ZEPYA,98,714;%%
  %879 citations counted in INSPIRE as of 06 May 2013

\bibitem{Weisskopf}
V.~Weisskopf, 
Kong.\ Dans.\ Vid.\ Selsk., Mat.-fys.\ Medd.\ {\bf XIV}, 6 (1936).

\bibitem{Toll:1952}
J.~S.~Toll,
Ph.D. thesis, Princeton Univ., 1952 (unpublished).

%\cite{Baier}
\bibitem{Baier}
R.~Baier and P.~Breitenlohner, 
{Act.~Phys.~Austriaca} {\bf 25}, 212 (1967); 
{Nuov.~Cim.~B}\ {\bf 47} 117 (1967).

%\cite{BialynickaBirula:1970vy}
\bibitem{BialynickaBirula:1970vy}
  Z.~Bialynicka-Birula and I.~Bialynicki-Birula,
  %``Nonlinear effects in Quantum Electrodynamics. Photon propagation and photon splitting in an external field,''
  Phys.\ Rev.\ D {\bf 2}, 2341 (1970).
  %doi:10.1103/PhysRevD.2.2341
  %%CITATION = doi:10.1103/PhysRevD.2.2341;%%
  %225 citations counted in INSPIRE as of 08 Aug 2017
  
%\cite{Heinzl:2006xc}
\bibitem{Heinzl:2006xc} 
  T.~Heinzl, {\it et al.}, 
  %B.~Liesfeld, K.~U.~Amthor, H.~Schwoerer, R.~Sauerbrey and A.~Wipf,
  %``On the observation of vacuum birefringence,''
  Opt.\ Commun.\  {\bf 267}, 318 (2006).
  %[hep-ph/0601076].
  %%CITATION = HEP-PH/0601076;%%
  %92 citations counted in INSPIRE as of 16 Apr 2015

%\cite{DiPiazza:2006pr}
\bibitem{DiPiazza:2006pr} 
  A.~Di Piazza, K.~Z.~Hatsagortsyan and C.~H.~Keitel,
  %``Light diffraction by a strong standing electromagnetic wave,''
  Phys.\ Rev.\ Lett.\  {\bf 97}, 083603 (2006)
  [hep-ph/0602039].
  %%CITATION = HEP-PH/0602039;%%
  %53 citations counted in INSPIRE as of 01 Dec 2014

%\cite{Dinu:2013gaa}
\bibitem{Dinu:2013gaa} 
  V.~Dinu, T.~Heinzl, A.~Ilderton, M.~Marklund and G.~Torgrimsson,
  %``Vacuum refractive indices and helicity flip in strong-field QED,''
  Phys.\ Rev.\ D {\bf 89}, 125003 (2014)
  [arXiv:1312.6419 [hep-ph]]; %.
  %%CITATION = ARXIV:1312.6419;%%
  %7 citations counted in INSPIRE as of 01 Dec 2014
%\cite{Dinu:2014tsa}
%\bibitem{Dinu:2014tsa} 
%  V.~Dinu, T.~Heinzl, A.~Ilderton, M.~Marklund and G.~Torgrimsson,
  %``Photon polarisation in light-by-light scattering: finite size effects,''
  Phys.\ Rev.\ D {\bf 90}, 045025 (2014)
  [arXiv:1405.7291 [hep-ph]].
  %%CITATION = ARXIV:1405.7291;%%
  %6 citations counted in INSPIRE as of 01 Dec 2014
 
%\cite{Karbstein:2015xra}
\bibitem{Karbstein:2015xra} 
  F.~Karbstein, H.~Gies, M.~Reuter and M.~Zepf,
  %``Vacuum birefringence in strong inhomogeneous electromagnetic fields,''
  Phys.\ Rev.\ D {\bf 92}, 071301 (2015)
  [arXiv:1507.01084 [hep-ph]].
  %%CITATION = ARXIV:1507.01084;%%
  %1 citations counted in INSPIRE as of 02 Nov 2015
  
%\cite{Schlenvoigt:2016}
\bibitem{Schlenvoigt:2016} 
 H.~-P.~Schlenvoigt, T.~Heinzl, U.~Schramm, T.~Cowan and R.~Sauerbrey,
 %``Detecting vacuum birefringence with x-ray free electron lasers and high-power optical lasers: a feasibility study,''
 Physica\ Scripta {\bf 91}, 023010 (2016).
  
%\cite{Karbstein:2016lby}
\bibitem{Karbstein:2016lby} 
  F.~Karbstein and C.~Sundqvist,
  %``Probing vacuum birefringence using x-ray free electron and optical high-intensity lasers,''
  Phys.\ Rev.\ D {\bf 94}, 013004 (2016)
  %doi:10.1103/PhysRevD.94.013004
  [arXiv:1605.09294 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.94.013004;%%
  %5 citations counted in INSPIRE as of 24 Sep 2017

%\cite{Kotkin:1996nf}
\bibitem{Kotkin:1996nf} 
  G.~L.~Kotkin and V.~G.~Serbo,
  %``Variation in polarization of high-energy gamma quanta traversing a bunch of polarized laser photons,''
  Phys.\ Lett.\ B {\bf 413}, 122 (1997).
  %[hep-ph/9611345].
  %%CITATION = HEP-PH/9611345;%%
  %24 citations counted in INSPIRE as of 29 Jun 2015
  
%\cite{Nakamiya:2015pde}
\bibitem{Nakamiya:2015pde} 
  Y.~Nakamiya, K.~Homma, T.~Moritaka and K.~Seto,
  %``Probing vacuum birefringence under a high-intensity laser field with gamma-ray polarimetry at the GeV scale,''
  Phys.\ Rev.\ D {\bf 96}, 053002 (2017)
  %doi:10.1103/PhysRevD.96.053002
  [arXiv:1512.00636 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.96.053002;%%
  %4 citations counted in INSPIRE as of 21 Oct 2017
  
%\cite{Ilderton:2016khs}
\bibitem{Ilderton:2016khs} 
  A.~Ilderton and M.~Marklund,
  %``Prospects for studying vacuum polarisation using dipole and synchrotron radiation,''
  J.\ Plasma Phys.\  {\bf 82}, 655820201 (2016)
  %doi:10.1017/S0022377816000192
  [arXiv:1601.08045 [hep-ph]].
  %%CITATION = doi:10.1017/S0022377816000192;%%
  %4 citations counted in INSPIRE as of 29 Sep 2017

%\cite{King:2016jnl}
\bibitem{King:2016jnl} 
  B.~King and N.~Elkina,
  %``Vacuum birefringence in high-energy laser-electron collisions,''
  Phys.\ Rev.\ A {\bf 94}, 062102 (2016)
  %doi:10.1103/PhysRevA.94.062102
  [arXiv:1603.06946 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevA.94.062102;%%
  %4 citations counted in INSPIRE as of 29 Sep 2017
 
%\cite{Bragin:2017yau}
\bibitem{Bragin:2017yau} 
  S.~Bragin, S.~Meuren, C.~H.~Keitel and A.~Di Piazza,
  %``High-energy vacuum birefringence and dichroism in an ultrastrong laser field,''
  arXiv:1704.05234 [hep-ph].
  %%CITATION = ARXIV:1704.05234;%%

\bibitem{Hatsagortsyan:2011}
K.~Z.~Hatsagortsyan and G.~Y.~Kryuchkyan,
%``Bragg Scattering of Light in Vacuum Structured by Strong Periodic Fields,''
 Phys.\ Rev.\ Lett. {\bf 107}, 053604 (2011).
  
%\cite{King:2013am}
 \bibitem{King:2013am} 
  B.~King, A.~Di Piazza and C.~H.~Keitel,
  %``A matterless double slit,''
 Nature Photon.\ {\bf 4}, 92 (2010) 
  [arXiv:1301.7038 [physics.optics]]; %.
  %%CITATION = ARXIV:1301.7038;%%
%\cite{King:2013zz}
%\bibitem{King:2013zz} 
%  B.~King, A.~Di Piazza and C.~H.~Keitel,
  %``Double-slit vacuum polarisation effects in ultra-intense laser fields,''
Phys.\ Rev.\ A {\bf 82}, 032114 (2010)
  [arXiv:1301.7008 [physics.optics]].
  %%CITATION = ARXIV:1301.7008;%%

%\cite{Tommasini:2010fb}
\bibitem{Tommasini:2010fb} 
  D.~Tommasini and H.~Michinel,
  %``Light by light diffraction in vacuum,''
  Phys.\ Rev.\ A {\bf 82}, 011803 (2010)
  [arXiv:1003.5932 [hep-ph]].
  %%CITATION = ARXIV:1003.5932;%%
  %6 citations counted in INSPIRE as of 06 May 2013

%\cite{King:2012aw}
\bibitem{King:2012aw} 
  B.~King and C.~H.~Keitel,
  %``Photon-photon scattering in collisions of laser pulses,''
  New J.\ Phys.\  {\bf 14}, 103002 (2012)
  [arXiv:1202.3339 [hep-ph]].
  %%CITATION = ARXIV:1202.3339;%%

%\cite{Gies:2013yxa}
\bibitem{Gies:2013yxa} 
  H.~Gies, F.~Karbstein and N.~Seegert,
  %``Quantum Reflection as a New Signature of Quantum Vacuum Nonlinearity,''
  New J.\ Phys.\  {\bf 15}, 083002 (2013)
  %doi:10.1088/1367-2630/15/8/083002
  [arXiv:1305.2320 [hep-ph]];
  %%CITATION = doi:10.1088/1367-2630/15/8/083002;%%
  %24 citations counted in INSPIRE as of 29 Sep 2017
%\cite{Gies:2014wsa}
%\bibitem{Gies:2014wsa} 
%  H.~Gies, F.~Karbstein and N.~Seegert,
  %``Quantum reflection of photons off spatio-temporal electromagnetic field inhomogeneities,''
  New J.\ Phys.\  {\bf 17}, 043060 (2015)
  %doi:10.1088/1367-2630/17/4/043060
  [arXiv:1412.0951 [hep-ph]].
  %%CITATION = doi:10.1088/1367-2630/17/4/043060;%%
  %11 citations counted in INSPIRE as of 29 Sep 2017

%\cite{Dittrich:2000zu}
\bibitem{Dittrich:2000zu} 
  W.~Dittrich and H.~Gies,
  %``Probing the quantum vacuum. Perturbative effective action approach in quantum electrodynamics and its application,''
  Springer Tracts Mod.\ Phys.\  {\bf 166}, 1 (2000).
  %%CITATION = STPHB,166,1;%%
  %130 citations counted in INSPIRE as of 28 Oct 2015
  
%\cite{Marklund:2008gj}
\bibitem{Marklund:2008gj} 
  M.~Marklund and J.~Lundin,
  %``Quantum Vacuum Experiments Using High Intensity Lasers,''
  Eur.\ Phys.\ J.\ D {\bf 55}, 319 (2009)
  [arXiv:0812.3087 [hep-th]].
  %%CITATION = ARXIV:0812.3087;%%
  %18 citations counted in INSPIRE as of 06 May 2013

%\cite{DiPiazza:2011tq}
\bibitem{DiPiazza:2011tq} 
  A.~Di Piazza, C.~M\"uller, K.~Z.~Hatsagortsyan and C.~H.~Keitel,
  %``Extremely high-intensity laser interactions with fundamental quantum systems,''
  Rev.\ Mod.\ Phys.\  {\bf 84}, 1177 (2012)
  [arXiv:1111.3886 [hep-ph]].
  %%CITATION = ARXIV:1111.3886;%%
  %34 citations counted in INSPIRE as of 06 May 2013

%\cite{Battesti:2012hf}
\bibitem{Battesti:2012hf} 
  R.~Battesti and C.~Rizzo,
  %``Magnetic and electric properties of quantum vacuum,''
  Rept.\ Prog.\ Phys.\  {\bf 76}, 016401 (2013)
  [arXiv:1211.1933 [physics.optics]].
  %%CITATION = ARXIV:1211.1933;%%
  %13 citations counted in INSPIRE as of 01 Dec 2014

%\cite{King:2015tba}
\bibitem{King:2015tba} 
  B.~King and T.~Heinzl,
  %``Measuring Vacuum Polarisation with High Power Lasers,''
  High Power Laser Science and Engineering, 4, e5 (2016)
  %doi:10.1017/hpl.2016.1
  [arXiv:1510.08456 [hep-ph]].
  %%CITATION = doi:10.1017/hpl.2016.1;%%
  %15 citations counted in INSPIRE as of 11 Jul 2017  
%\cite{Cantatore:2008zz}
\bibitem{Cantatore:2008zz} 
  G.~Cantatore [PVLAS Collaboration],
  %``Recent results from the PVLAS experiment on the magnetized vacuum,''
  Lect.\ Notes Phys.\  {\bf 741}, 157 (2008); %.
  %%CITATION = LNPHA,741,157;%%
  %3 citations counted in INSPIRE as of 06 May 2013
%\cite{Zavattini:2007ee}
%\bibitem{Zavattini:2007ee} 
  E.~Zavattini {\it et al.} [PVLAS Collaboration],
  %``New PVLAS results and limits on magnetically induced optical rotation and ellipticity in vacuum,''
  Phys.\ Rev.\ D {\bf 77}, 032006 (2008);
  %[arXiv:0706.3419 [hep-ex]]; %.
  %%CITATION = ARXIV:0706.3419;%%
  %130 citations counted in INSPIRE as of 06 May 2013
%\cite{DellaValle:2013xs}
%\bibitem{DellaValle:2013xs} 
  F.~Della Valle {\it et al.},
  %``Measurements of vacuum magnetic birefringence using permanent dipole magnets: the PVLAS experiment,''
  New\ J.\ Phys.\ {\bf 15} 053026 (2013);
  %arXiv:1301.4918 [quant-ph].
  %%CITATION = ARXIV:1301.4918;%%
%\cite{DellaValle:2015xxa}
%\bibitem{DellaValle:2015xxa} 
  F.~Della Valle,  {\it et al.},
  %``The PVLAS experiment: measuring vacuum magnetic birefringence and dichroism with a birefringent Fabry–Perot cavity,''
  Eur.\ Phys.\ J.\ C {\bf 76}, 24 (2016).
  %doi:10.1140/epjc/s10052-015-3869-8
  %[arXiv:1510.08052 [physics.optics]].
  %%CITATION = doi:10.1140/epjc/s10052-015-3869-8;%%
  %2 citations counted in INSPIRE as of 25 Feb 2016

%\cite{Berceau:2011zz}
\bibitem{Berceau:2011zz} 
  P.~Berceau, {\it et al.},
  %R.~Battesti, M.~Fouche and C.~Rizzo,
  %``The vacuum magnetic birefringence experiment: A test for quantum electrodynamics,''
  Can.\ J.\ Phys.\  {\bf 89}, 153 (2011);
  %%CITATION = CJPHA,89,153;%%
  P.~Berceau, {\it et al.},
  %M.~Fouche, R.~Battesti and C.~Rizzo,
  %``Magnetic linear birefringence measurements using pulsed fields,''
  Phys.\ Rev.\ A, {\bf 85}, 013837 (2012);
  %[arXiv:1109.4792 [physics.optics]];
  A.~Cadene, {\it et al.},
  %P.~Berceau, M.~Fouche, R.~Battesti and C.~Rizzo,
  %``Vacuum magnetic linear birefringence using pulsed fields: status of the BMV experiment,''
  Eur.\ Phys.\ J.\ D {\bf 68}, 16 (2014).
  %[arXiv:1302.5389 [physics.optics]].
  %%CITATION = ARXIV:1302.5389;%%

%\cite{Fan:2017fnd}
\bibitem{Fan:2017fnd} 
  X.~Fan {\it et al.},
  %``The OVAL experiment: A new experiment to measure vacuum magnetic birefringence using high repetition pulsed magnets,''
  Eur.\ Phys.\ J.\ D {\bf 71}, 308 (2017)
  %doi:10.1140/epjd/e2017-80290-7
  [arXiv:1705.00495 [physics.ins-det]].
  %%CITATION = doi:10.1140/epjd/e2017-80290-7;%%
  %4 citations counted in INSPIRE as of 01 Dec 2017

%\cite{Inada:2017lop}
\bibitem{Inada:2017lop} 
  T.~Inada, {\it et al.},
  %T.~Yamazaki, T.~Yamaji, Y.~Seino, X.~Fan, S.~Kamioka, T.~Namba and S.~Asai,
  %``Probing Physics in Vacuum Using an X-ray Free-Electron Laser, a High-Power Laser, and a High-Field Magnet,''
  Appl. Sci. {\bf 7}, 671 (2017)
  %doi:10.3390/app7070671
  [arXiv:1707.00253 [hep-ex]].
  %%CITATION = doi:10.3390/app7070671;%%

\bibitem{narozhnyi:1968}
  N.~B.~Narozhnyi,
 %``Propagation of plane electromagnetic waves in a constant field,''
 Zh.\ Eksp.\ Teor.\ Fiz.\ {\bf 55}, 714 (1968)
 [Sov.\ Phys.\ JETP \textbf{28}, 371 (1969)].

\bibitem{BatShab}
  I.~A.~Batalin and A.~E.~Shabad,
  Zh.\ Eksp.\ Teor.\ Fiz.\  {\bf 60}, 894 (1971)
  [Sov.\ Phys.\ JETP\ {\bf 33}, 483 (1971)].

%\cite{Baier:1975ff}
\bibitem{Baier:1975ff} 
  V.~N.~Baier, A.~I.~Milshtein and V.~M.~Strakhovenko,
  %``Interaction Between a Photon and a High Intensity Electromagnetic Wave,''
  Zh.\ Eksp.\ Teor.\ Fiz.\  {\bf 69}, 1893 (1975)
  [Sov.\ Phys.\ JETP\ {\bf 42}, 961 (1976)].
  %%CITATION = ZETFA,69,1893;%%

%\cite{Becker:1974en}
\bibitem{Becker:1974en} 
  W.~Becker and H.~Mitter,
  %``Vacuum Polarization in Laser Fields,''
   J.\ Phys.\ A:\ Math.\ Gen.\ {\bf 8} 1638 (1975).
  %Print-74-0875 (TUBINGEN).
  %%CITATION = PRINT-74-0875 (TUBINGEN);%%  

%\cite{Karbstein:2015cpa}
\bibitem{Karbstein:2015cpa} 
  F.~Karbstein and R.~Shaisultanov,
  %``Photon propagation in slowly varying inhomogeneous electromagnetic fields,''
  Phys.\ Rev.\ D {\bf 91}, 085027 (2015)
  %doi:10.1103/PhysRevD.91.085027
  [arXiv:1503.00532 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.91.085027;%%
  %12 citations counted in INSPIRE as of 08 Aug 2017

%\cite{Gusynin:1995bc}
\bibitem{Gusynin:1995bc} 
  V.~P.~Gusynin and I.~A.~Shovkovy,
  %``Derivative expansion for the one loop effective Lagrangian in QED,''
  Can.\ J.\ Phys.\  {\bf 74}, 282 (1996)
  %doi:10.1139/p96-044
  [hep-ph/9509383];
  %%CITATION = doi:10.1139/p96-044;%%
  %54 citations counted in INSPIRE as of 24 Mar 2017
%\cite{Gusynin:1998bt}
%\bibitem{Gusynin:1998bt} 
%  V.~P.~Gusynin and I.~A.~Shovkovy,
  %``Derivative expansion of the effective action for QED in (2+1)-dimensions and (3+1)-dimensions,''
  J.\ Math.\ Phys.\  {\bf 40}, 5406 (1999)
  %doi:10.1063/1.533037
  [hep-th/9804143].
  %%CITATION = doi:10.1063/1.533037;%%
  %63 citations counted in INSPIRE as of 24 Mar 2017
   
%\cite{Dunne:2000up}
\bibitem{Dunne:2000up} 
  G.~V.~Dunne,
  %``QED effective actions in inhomogeneous backgrounds: Summing the derivative expansion,''
  AIP Conf.\ Proc.\  {\bf 564}, 247 (2001)
  %doi:10.1063/1.1374991
  [hep-th/0011036].
  %%CITATION = doi:10.1063/1.1374991;%%
  %4 citations counted in INSPIRE as of 16 Mar 2017

%\cite{Karbstein:2017jgh}
\bibitem{Karbstein:2017jgh} 
  F.~Karbstein and E.~A.~Mosman,
  %``The photon polarization tensor in pulsed Hermite- and Laguerre-Gaussian beams,''
  Phys.\ Rev.\ D {\bf 96}, 116004 (2017)
  %doi:10.1103/PhysRevD.96.116004
  [arXiv:1711.06151 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.96.116004;%%

%\cite{Galtsov:1982}
\bibitem{Galtsov:1982} 
  G.~V.~Galtsov and N.~S.~Nikitina,
  %``Macroscopic vacuum effects in an inhomogeneous and nonstationary electromagnetic field,''
  Zh.\ Eksp.\ Teor.\ Fiz.\  {\bf 84}, 1217 (1983) [Sov.\ Phys.\ JETP\ {\bf 57}, 705 (1983)].

%\cite{Davis:1979}
\bibitem{Davis:1979} 
  L.~W.~Davis,
  %``Theory of electromagnetic beams,''
  Phys.\ Rev.\ A {\bf 19}, 3 (1979).
  
%\cite{Barton:1989}
\bibitem{Barton:1989}
 J.~P.~Barton and D.~R.~Alexander,
 %``Fifth‐order corrected electromagnetic field components for a fundamental Gaussian beam,''
 J. Appl. Phy. {\bf 66}, 2800 (1989).
 
%\cite{Salamin:2006ff}
\bibitem{Salamin:2006ff} 
  Y.~I.~Salamin, S.~X.~Hu, K.~Z.~Hatsagortsyan and C.~H.~Keitel,
  %``Relativistic high-power laser-matter interactions,''
  Phys.\ Rept.\  {\bf 427}, 41 (2006).
  %%CITATION = PRPLC,427,41;%%
  %88 citations counted in INSPIRE as of 20 Feb 2015

\bibitem{Siegman}
A.~E.~Siegman, \textit{Lasers}, 1st. ed. (University Science Books, Herndon, VA, 1986)

\end{thebibliography}

\documentclass[a4paper]{jpconf}
\usepackage{ae}
\usepackage{bm}
\usepackage{color}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{slashed}
\usepackage{wasysym}
\usepackage{setspace}
\usepackage{subfigure}
\usepackage{ulem}
\usepackage{cite}

\newcommand{\lambdabar}{{\mkern0.75mu\mathchar '26\mkern -9.75mu\lambda}}

\newcommand{\Eqref}[1]{Eq.~\eqref{#1}}

\begin{document}

\title{Photon polarization tensor in circularly polarized Hermite- and Laguerre-Gaussian beams}

\author{Felix Karbstein$^{1,2}$ and Elena A. Mosman$^{3,4}$}

\address{$^1$ Helmholtz-Institut Jena, Fr\"obelstieg 3, 07743 Jena, Germany} 

\address{$^2$ Theoretisch-Physikalisches Institut, Abbe Center of Photonics, Friedrich-Schiller-Universit\"at Jena, Max-Wien-Platz 1, 07743 Jena, Germany}

\address{$^3$ National Research Tomsk Polytechnic University, Lenin Ave. 30, 634050 Tomsk, Russia}

\address{$^4$ Physics Faculty, Tomsk State University, Lenin Ave. 36, 634050 Tomsk, Russia}

\ead{felix.karbstein@uni-jena.de, mosmanea@tpu.ru}

\begin{abstract}
 We derive analytical expressions for the photon polarization tensor in circularly polarized Hermite- and Laguerre-Gaussian beams, complementing the corresponding results for linearly polarized beams obtained recently.
 As they are based upon a locally constant field approximation of the one-loop Heisenberg-Euler effective Lagrangian for quantum electrodynamics (QED) in constant fields, our results are generically limited to slowly varying electromagnetic fields, varying on spatial (temporal) scales much larger than the Compton wavelength (time) of the electron. 
\end{abstract}

\section{Introduction}

The photon polarization tensor is a central object in the study of probe photon propagation in the quantum vacuum.
It encodes quantum corrections to the classical propagation of light.
In the absence of electromagnetic fields, by construction quantum fluctuations do not change the vacuum speed of light and do not result in any modifications for photon propagation.
This is different in external fields: As electromagnetic fields couple to charges, they can {\it polarize} virtual charged particle-antiparticle fluctuations \cite{Heisenberg:1935qt,Weisskopf}, giving rise to remarkable effects, such as vacuum birefringence \cite{Toll:1952,Baier,BialynickaBirula:1970vy,Heinzl:2006xc,DiPiazza:2006pr,Dinu:2013gaa,Karbstein:2015xra,Schlenvoigt:2016,Karbstein:2016lby,Kotkin:1996nf,Nakamiya:2015pde,Ilderton:2016khs,King:2016jnl,Bragin:2017yau} and -- in inhomogeneous fields -- vacuum diffraction effects \cite{Hatsagortsyan:2011,King:2013am,Tommasini:2010fb,King:2012aw,Gies:2013yxa} experienced by probe photons; cf. also the reviews \cite{Dittrich:2000zu,Marklund:2008gj,DiPiazza:2011tq,Battesti:2012hf,King:2015tba} and references therein. 
For ongoing experimental efforts aiming at the detection of these effects, see Refs.~\cite{Cantatore:2008zz,Berceau:2011zz,Fan:2017fnd,Inada:2017lop}.
Exact analytical results for the photon polarization tensor $\Pi^{\rho\sigma}$ in external field QED
at one-loop level $\sim\alpha/\pi$, where\footnote{Throughout this article, we use the Heaviside-Lorentz system and units where $c=\hbar=1$. Our metric convention is $g^{\mu\nu}={\rm diag}(-1,+1,+1,+1)$.} $\alpha=e^2/(4\pi)\simeq1/137$ is the fine-structure constant, with elementary charge $e$, are available for constant electromagnetic fields \cite{BatShab,narozhnyi:1968} and plane wave backgrounds \cite{Baier:1975ff,Becker:1974en}.

Besides, approximate results for $\Pi^{\rho\sigma}$ in slowly varying inhomogeneous electromagnetic fields have been determined on the basis of a locally constant field approximation (LCFA) \cite{Karbstein:2015cpa} to the Heisenberg-Euler effective Lagrangian in constant fields \cite{Heisenberg:1935qt}.
For insights into the Heisenberg-Euler effective Lagrangian beyond constant fields, cf. the derivative expansion results \cite{Gusynin:1995bc,Dunne:2000up}.
The LCFA allows for reliable insights in electromagnetic field configurations which vary on spatial (temporal) scales much larger than the Compton wavelength $\lambdabar=1/m_e\approx3.86\cdot10^{-13}\,{\rm m}$ (time $\lambdabar/c\approx1.29\cdot10^{-21}\,{\rm s}$) of the electron; $m_e\approx511\,{\rm keV}$ is the electron mass.
So far, this approach has been adopted to determine $\Pi^{\rho\sigma}$ for linearly polarized paraxial Hermite- and Laguerre-Gaussian beams of arbitrary mode composition \cite{Karbstein:2017jgh}.

In these proceedings, we provide the analogous expressions for circular polarized beams.

\section{Results}

\subsection{Field configuration}

The electric and magnetic field vectors describing a generic superposition of co-propagating, electromagnetic waves in vacuum can be represented as 
\begin{equation}
 \vec{E}=\sum_{\cal N}{\cal E}_{\cal N}\bigl(\hat{\vec{e}}_{\rm x}\cos\phi_{\cal N}+\hat{\vec{e}}_{\rm y}\sin\phi_{{\cal N}}\bigr)\,,\quad
 \vec{B}=\vec{E}\big|_{\phi_{\cal N}\to\phi_{\cal N}+\frac{\pi}{2}}\,, \label{eq:Fields}
\end{equation}
where ${\cal E}_{\cal N}$ denote amplitude profile and $\phi_{\cal N}$ a phase function of mode $\cal N$, respectively.
Without loss of generality, in \Eqref{eq:Fields} we assumed the waves to propagate along the positive $\rm z$ axis.
Aiming at the determination of $\Pi^{\rho\sigma}$ in this field configuration, it is helpful to note that \Eqref{eq:Fields} implies
\begin{align}
    {\cal F}&=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}
    =0\,,\nonumber \\
    {\cal G}&=\frac{1}{4}F_{\mu\nu}{}^*\!F^{\mu\nu}=-\sum_{\cal N}\sum_{{\cal N}'}{\cal E}_{\cal N}{\cal E}_{{\cal N}'}\sin(\phi_{\cal N}-\phi_{{\cal N}'}) , \nonumber\\
    (k F)^\rho&=\sum_{\cal N}{\cal E}_{\cal N}\bigl[\varepsilon_2^\rho(k)\sin\phi_{\cal N}+\varepsilon_1^\rho(k)\cos\phi_{\cal N}\bigr], \nonumber \\
    (k{}^*\!F)^\rho&=\sum_{\cal N}{\cal E}_{\cal N}\bigl[\varepsilon_2^\rho(k)\cos\phi_{\cal N}-\varepsilon_1^\rho(k)\sin\phi_{\cal N}\bigr], \label{eq:bb}
\end{align}
with $k^\mu=(\omega,\vec{k})$, $(kF)^\mu=k_\nu F^{\nu\mu}$
and dual field strength tensor ${}^*\!F^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\alpha\beta}F_{\alpha\beta}$.
The four-vectors $\varepsilon_{1,2}^\mu(k)$ are defined as \cite{Karbstein:2015cpa}
\begin{equation}
 \varepsilon_1^\mu(k):=(-k_{\rm x},k_{\rm z}-\omega,0,-k_{\rm x})
 \,,\quad
 \varepsilon_2^\mu(k):=(-k_{\rm y},0,k_{\rm z}-\omega,-k_{\rm y})\,.
\end{equation}

\subsection{Photon polarization tensor}

Resorting to a LCFA, the one-loop photon polarization tensor in the background field~\eqref{eq:Fields} can be expressed as \cite{Karbstein:2015cpa}\footnote{For completeness, note that this definition of $\Pi^{\rho\sigma}$ differs from that of Ref.~\cite{Karbstein:2015cpa} by an overall minus sign.}
\begin{equation}
 \Pi^{\rho\sigma}(k,k')
 =-\frac{\alpha}{\pi}\frac{1}{45}\Bigl(\frac{e}{m_e^2}\Bigr)^2\int{\rm d}^4x\,{\rm e}^{{\rm i}(k+k')x} \Bigl[4 (kF)^\rho  (k'F)^\sigma
 + 7 (k{}^*\!F)^\rho (k'{}^*\!F)^\sigma+7{\cal G}\,
 k'_\mu k_\alpha \epsilon^{\rho\sigma\mu\alpha}
 \Bigr], \label{eq:Pi}
\end{equation}
where we neglected terms $\sim\upsilon^2\bigl[{\cal O}\bigl((\frac{\upsilon}{m_e})^2\bigr)+{\cal O}\bigl((\frac{eF}{m_e^2})^4\bigr)\bigr]$. Here, the momentum scale $\upsilon$ delimits the typical momentum scales of variation of the background fields as well as the photon energies $\omega,\omega'$ and momenta $|\vec{k}|,|\vec{k}'|$ from above \cite{Galtsov:1982,Karbstein:2015cpa}.
With the help of \Eqref{eq:bb}, \Eqref{eq:Pi} can be cast into the following form
\begin{align}
 \Pi^{\rho\sigma}(k,k') = -\frac{\alpha}{\pi}\frac{1}{45}\sum_{\ell=\pm1}
  \biggl\{&\!-\frac{7}{2}\,\Pi^-_\ell(k+k') (-{\rm i})^l k'_\mu k_\alpha\epsilon^{\rho\sigma\mu\alpha} \nonumber\\
  +\frac{11}{4}\,\Pi_\ell^-(k+k')&\Bigl[\varepsilon_1^\rho(k)\varepsilon_1^\sigma(k')+\varepsilon_2^\rho(k)\varepsilon_2^\sigma(k')
  +(-{\rm i})^l\bigl[\varepsilon_2^\rho(k)\varepsilon_1^\sigma(k')
 -\varepsilon_1^\rho(k)\varepsilon_2^\sigma(k')]\Bigr] \nonumber\\
  -\frac{3}{4}\,\Pi^+_\ell(k+k')&\Bigl[\varepsilon_1^\rho(k)\varepsilon_1^\sigma(k')-\varepsilon_2^\rho(k)\varepsilon_2^\sigma(k')
  +(-{\rm i})^l\bigl[\varepsilon_2^\rho(k)\varepsilon_1^\sigma(k')+\varepsilon_1^\rho(k)\varepsilon_2^\sigma(k')\bigr]\Bigr] \biggr\} , \label{eq:Pi2}
\end{align}
where we made use of the definitions
\begin{equation}
 \Pi_\ell^\pm(k+k') = \sum_{\cal N}\sum_{{\cal N}'}\,(\Pi_\ell^\pm)_{{\cal N};{\cal N}'}(k+k')\,, \label{eq:Pi2b}
\end{equation}
and
\begin{equation}
 (\Pi_\ell^\pm)_{{\cal N};{\cal N}'}(k+k')=\int{\rm d}^4x\,{\rm e}^{{\rm i}(k+k')x}\,\frac{e{\cal E}_{\cal N}}{m_e^2}\,\frac{e{\cal E}_{{\cal N}'}}{m_e^2}\,{\rm e}^{{\rm i}\ell(\phi_{\cal N}\pm\phi_{{\cal N}'})} \,. \label{eq:Pi2c}
\end{equation}

\subsection{Circular polarization}

The above results can be straightforwardly adopted to describe the photon polarization tensor in circularly polarized paraxial beams at leading order in the diffraction angle $\theta\ll1$ (cf. below) \cite{Davis:1979,Barton:1989,Salamin:2006ff}.
In this case, we have
\begin{align}
 {\cal E}_{\cal N}\quad&\to\quad{\cal E}_{\cal N}(x)\,, \nonumber\\
 \phi_{\cal N}\quad&\to\quad\pm\Phi_{\cal N}(x)\,,
\end{align}
where $+/-$ refers to right-/left-handed circular polarization. For the explicit mode amplitude ${\cal E}_{\cal N}(x)$ and phase profiles $\phi_{\cal N}(x)$ adopted in these proceedings, see below. 
Hence, in our explicit calculations we can without loss of generality stick to right-handed circular polarization, $\phi_{\cal N}\to\Phi_{\cal N}(x)$.
The result for left-handed circular polarization follows therefrom by substituting $\Pi_\ell^\pm(k+k')\to\Pi_{-\ell}^\pm(k+k')$ in \Eqref{eq:Pi2}.

\subsection{Field profiles}

Hermite-Gaussian (HG) modes are labeled by ${\cal N}=\{m\in\mathbb{N}_0,n\in\mathbb{N}_0\}$. They exhibit a Cartesian symmetry about the beam's propagation direction and are characterized by the following amplitude and phase profiles \cite{Siegman},
\begin{align}
 {\cal E}_{m,n}(x)&={\mathfrak E}_{m,n}\,H_m\Bigl(\tfrac{\sqrt{2}{\rm x}}{w({\rm z})}\Bigr)H_n\Bigl(\tfrac{\sqrt{2}{\rm y}}{w({\rm z})}\Bigr)\,{\rm e}^{-\frac{({\rm z}-t)^2}{(\tau/2)^2}}\,\frac{w_0}{w({\rm z})}\,{\rm e}^{-\frac{r^2}{w^2({\rm z})}}, \nonumber\\
  \Phi_{m,n}(x)&=
    \Omega({\rm z}-t)+\frac{\rm z}{{\rm z}_R}\frac{r^2}{w^2({\rm z})}-(m+n+1)\arctan\Bigl(\frac{\rm z}{{\rm z}_R}\Bigr)+\varphi_{n,m}\,, \label{eq:HG}
\end{align}
where $H_m(\chi)$ denote Hermite polynomials, $\Omega=2\pi/\lambda$ is the frequency of the beam, and $w({\rm z})=w_0\sqrt{1+({\rm z}/{\rm z}_R)^2}$, with waist size $w_0$ and Rayleigh range ${\rm z}_R=\pi w_0^2/\lambda$.
The constants ${\mathfrak E}_{\cal N}$ and $\varphi_{\cal N}$ are a mode specific peak field amplitude and phase offset, respectively, and $\tau$ is the pulse duration.

Analogously, Laguerre-Gaussian (LG) modes are labeled by ${\cal N}=\{l\in\mathbb{Z},p\in\mathbb{N}_0\}$. The transverse profiles of LG modes are circularly symmetric and, resorting to polar coordinates, $r=\sqrt{{\rm x}^2+{\rm y}^2}$ and $\varphi={\rm arg}({\rm x}+{\rm iy})$, read  \cite{Siegman}
\begin{align}
 {\cal E}_{l,p}(x)&={\mathfrak E}_{l,p} \biggl(\frac{\sqrt{2}r}{w({\rm z})}\biggr)^{|l|} L^{|l|}_p\Bigl(\bigl(\tfrac{\sqrt{2}r}{w({\rm z})}\bigr)^2\Bigr) \,{\rm e}^{-\frac{({\rm z}-t)^2}{(\tau/2)^2}}\,\frac{w_0}{w({\rm z})}\,{\rm e}^{-\frac{r^2}{w^2({\rm z})}}, \nonumber\\
  \Phi_{l,p}(x)&=
    \Omega({\rm z}-t)+\frac{\rm z}{{\rm z}_R}\frac{r^2}{w^2({\rm z})}-(|l|+2p+1)\arctan\Bigl(\frac{\rm z}{{\rm z}_R}\Bigr)-l\varphi+\varphi_{l,p}\,, \label{eq:LG}
\end{align}
where $L_p^{|l|}(\chi)$ are generalized Laguerre polynomials. For further details, cf. Ref.~\cite{Karbstein:2017jgh}.

For given laser parameters, i.e., total pulse energy $W$, pulse duration $\tau$ and waist radius of the fundamental Gaussian mode $w_0$, the total energy is partitioned into the different modes ${\cal N}$ as $W=\sum_{\cal N}W_{\cal N}$, where $W_{\cal N}$ denotes the energy put in mode $\cal N$.
Making use of the fact that the intensity associated with the electromagnetic field in mode $\cal N$ is given by $I_{\cal N}=\vec{E}_{\cal N}^2={\cal E}_{\cal N}^2$, where $\vec{E}_{\cal N}={\cal E}_{\cal N}(\hat{\vec{e}}_{\rm x}\cos\phi_{\cal N}+\hat{\vec{e}}_{\rm y}\sin\phi_{{\cal N}})$, it is straightforward to derive the following relation between $W_{\cal N}$ and the peak field amplitude $\mathfrak{E}_{\cal N}$,
\begin{equation}
W_{\cal N}=\Bigl(\frac{\pi}{2}\Bigr)^{\frac{3}{2}}\,{\mathfrak c}_{\cal N}\,{\mathfrak E}_{\cal N}^2\,\frac{\tau}{2}\,w_0^2
\quad\leftrightarrow\quad
{\mathfrak E}_{\cal N}^2=4\sqrt{\frac{2}{\pi}}\frac{1}{{\mathfrak c}_{\cal N}}\frac{W_{\cal N}}{\pi w_0^2\tau}\,,
\label{eq:Wpulse}
\end{equation}
with mode specific coefficients ${\mathfrak c}_{\cal N}$ given by $c_{\cal N}\to{\mathfrak c}_{m,n}=2^{m+n}\,m!\,n!$ for HG modes,
and $c_{\cal N}\to{\mathfrak c}_{p,l}=(p+|l|)!/p!$ for LG modes.
For the details of the calculation, cf. the appendix of Ref.~\cite{Karbstein:2017jgh}.
We emphasize that the peak field amplitude squared in \Eqref{eq:Wpulse} amounts to half the one obtained for linearly polarized beams in Ref.~\cite{Karbstein:2017jgh}. The reason for this is that, in contrast to linear polarization, for circular polarization the phase factor $\phi_{\cal N}$ drops out completely when determining the mode intensity.

Inserting Eqs.~\eqref{eq:HG} and \eqref{eq:LG} into Eqs.~\eqref{eq:Pi2b} and \Eqref{eq:Pi2c} and performing the Fourier integrations, we obtain the following results for $\Pi^{\rho\sigma}$ in pulsed, circularly polarized LG and HG beams.

\subsection{Explicit results}

Following Ref.~\cite{Karbstein:2017jgh}, we represent the scalar coefficients~\eqref{eq:Pi2c} of the photon polarization as
\begin{align}
 (\Pi_\ell^\pm)_{{\cal N};{\cal N}'}(k+k')= \frac{e{\mathfrak E}_{\cal N}}{m_e^2}\frac{e{\mathfrak E}_{{\cal N}'}}{m_e^2}\,(2{\rm z}_R \, \pi w_0^2)\,
 \frac{\tau}{2}\sqrt{\frac{\pi}{2}}\,{\rm e}^{-\frac{1}{8}(\frac{\tau}{2})^2[\omega+\omega'+\ell(1\pm1)\Omega]^2}\,({\cal I}_\ell^\pm)_{{\cal N};{\cal N}'}(k+k') \,, \label{eq:Pi2c2}
\end{align}
where we encode the nontrivial momentum dependencies in the $\tau$ independent functions $({\cal I}_\ell^\pm)_{{\cal N};{\cal N}'}(k+k')$; note that $\lim_{\tau\to\infty}\frac{\tau}{2}\sqrt{\frac{\pi}{2}}\,{\rm e}^{-\frac{1}{8}(\frac{\tau}{2})^2\phi^2} =  2\pi\,\delta(\phi)$.
We will moreover make use of the definition
\begin{align}
 F_\Lambda\bigl(|a|,b\bigr):&=\int_{-\infty}^\infty\frac{\rm dz}{{\rm z}_R}\,
 \Bigl(\frac{w_0}{w({\rm z})}\Bigr)^\Lambda \,{\rm e}^{-{\rm i}a\frac{\rm z}{{\rm z}_R}-b\,(\frac{w({\rm z})}{w_0})^2} \nonumber\\
 &= \delta_{0,\Lambda}\sqrt{\frac{\pi}{b}}\,{\rm e}^{-\frac{1}{b}(\frac{a}{2})^2-b} + (1-\delta_{0,\Lambda}) \frac{\sqrt{\pi}}{\Gamma(\frac{\Lambda}{2})}\int_0^\infty\frac{{\rm d}s}{s}\,\frac{s^{\frac{\Lambda}{2}}}{\sqrt{s+b}}\,{\rm e}^{-\frac{1}{s+b}(\frac{a}{2})^2-(s+b)} \,, \label{eq:F}
\end{align}
for $b\geq0$, where $\delta_{0,\Lambda}$ is the Kronecker delta.
Equation~\eqref{eq:F} fulfills \cite{Karbstein:2017jgh},
\begin{equation}
 F_\Lambda\bigl(|a|,0\bigr)
 = \delta_{0,\Lambda}\, 2\pi\delta(a) + (1-\delta_{0,\Lambda})\, \frac{2\sqrt{\pi}}{\Gamma(\frac{\Lambda}{2})}\Bigl(\frac{|a|}{2}\Bigr)^{\frac{\Lambda-1}{2}}{\rm K}_{\frac{\Lambda-1}{2}}(|a|)\,,
 \label{eq:Fs}
\end{equation}
written in terms of the  Gamma function and the modified Bessel function of the second kind.
Employing the series representations of the Hermite polynomials and the generalized Laguerre polynomials as in Ref.~\cite{Karbstein:2017jgh}, the Fourier integrations can be performed explicitly and written in terms of $F_\Lambda\bigl(|a|,b\bigr)$ and parameter differentiations thereof.

For {\bf Hermite-Gaussian beams} we have ${\cal N}=\{m,n\}$, $N=m+n$, and obtain
\begin{align}
 ({\cal I}^\pm_\ell)_{{\cal N};{\cal N}'}^{{\rm HG}}(k)
 &=\frac{1}{4}\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{q=0}^{\lfloor\frac{n}{2}\rfloor}\sum_{j'=0}^{\lfloor\frac{m'}{2}\rfloor}\sum_{q'=0}^{\lfloor\frac{n'}{2}\rfloor}\frac{m!n!m'!n'!}{j!q!j'!q'!}
 \frac{(-{\rm i})^{N+N'}\, 2^{\frac{3}{2}(N+N')-3(j+j'+q+q')}}{(m-2j)!(n-2q)!(m'-2j')!(n'-2q')!} \nonumber\\
 &\hspace{7cm}\times ({\cal J}^\pm_\ell)_{{\cal N},j,q;{\cal N}',j',q'}^{{\rm HG}}(k) \,,
\end{align}
with
\begin{align}
 \!\!({\cal J}_\ell^+)_{{\cal N},j,q;{\cal N}',j',q'}^{{\rm HG}}(k)
 &={\rm e}^{{\rm i}\ell(\varphi_{m,n}+\varphi_{m',n'})}\bigl(\partial_{h_{\rm x}}\bigr)^{m+m'-2(j+j')}\,\bigl(\partial_{h_{\rm y}}\bigr)^{n+n'-2(q+q')} \nonumber\\
 &\quad\times\bigl(1+\ell\partial_{h_{\rm z}}\bigr)^{N+N'+1}\, {\rm e}^{-\frac{(w_0\vec{k}_\perp+\vec{h}_\perp)^2}{8}} \nonumber\\
 &\quad\times F_{2(N+N'+1-j-j'-q-q')}\!\bigl(\bigl|h_{\rm z}-{\rm z}_R(k\hat\kappa)+\ell\tfrac{(w_0\vec{k}_\perp+\vec{h}_\perp)^2}{8}\bigr|,0\bigr)\Big|_{\vec{h}=0} \,, \label{eq:IHGp}
\end{align}
and
\begin{align}
 ({\cal J}_\ell^-)_{{\cal N},j,q;{\cal N}',j',q'}^{{\rm HG}}(k)
 &= {\rm e}^{{\rm i} \ell(\varphi_{m,n}-\varphi_{m',n'})}\bigl[1+{\rm sign}\bigl(\ell(N-N')\bigr)\partial_{h_{\rm z}}\bigr]^{|N-N'|}  \bigl(\partial_{h_{\rm x}}\bigr)^{ 2\{\frac{m+m'}{2}\} }\nonumber \\
 &\quad\times\bigl(\partial_{h_{\rm y}}\bigr)^{2\{\frac{n+n'}{2}\}}\,\Bigl(\frac{1}{2}\partial_{c_{\rm x}}\Bigr)^{\lfloor\frac{m+m'}{2}\rfloor-j-j'}\,\Bigl(\frac{1}{2}\partial_{c_{\rm y}}\Bigr)^{\lfloor\frac{n+n'}{2}\rfloor-q-q'}\frac{1}{\sqrt{c_{\rm x}c_{\rm y}}} \nonumber\\
 &\quad\times F_{|N-N'|+2\{\frac{m+m'}{2}\}+2\{\frac{n+n'}{2}\}}\bigl(|h_{\rm z}-{\rm z}_R(k\hat\kappa)|,\Sigma_{i=1}^2\tfrac{(w_0k_i+h_i)^2}{8c_i}\bigr)\Big|_{\vec{c}=1,\,\vec{h}=0}\,. \label{eq:IHG0}
\end{align}
Here, we made use of the shorthand notations $\vec{k}_\perp=k_{\rm x}\vec{e}_{\rm x}+k_{\rm y}\vec{e}_{\rm y}$ and $(k\hat\kappa)=k_{\rm z}-\omega$ introduced in Ref.~\cite{Karbstein:2017jgh};
${\rm sign}(.)$ is the sign function. Besides, $\lfloor n\rfloor$ is the floor function which gives as output the largest integer less than or equal to $n$, and $2\{\frac{n}{2}\}:=n-2\lfloor\frac{n}{2}\rfloor$.

Conversely, for {\bf Laguerre-Gaussian beams} we have ${\cal N}=\{l,p\}$, $N=|l|+2p$, and obtain
\begin{equation}
 ({\cal I}^\pm_\ell)_{{\cal N};{\cal N}'}^{{\rm LG}}(k)
 =\frac{1}{4}\sum_{j=0}^p\sum_{j'=0}^{p'}\frac{(-\sqrt{2})^{|l|+|l'|}}{ j!j'!}\binom{p+|l|}{p-j}\binom{p'+|l'|}{p'-j'} \,
 ({\cal J}^\pm_\ell)_{{\cal N},j;{\cal N}',j'}^{{\rm LG}}(k)\,,
\end{equation}
with
\begin{align}
 ({\cal J}^+_\ell)_{{\cal N},j;{\cal N}',j'}^{{\rm LG}}(k)&={\rm e}^{{\rm i}\ell(\varphi_{l,p}+\varphi_{l',p'})}\,2^{j+j'}
  \bigl(1+\ell\partial_{h_{\rm z}}\bigr)^{N+N'+1}\bigl(\partial_{h_{\rm x}}^2+\partial_{h_{\rm y}}^2\bigr)^{j+j'} \nonumber\\
  &\quad\times\bigl({\rm i}\partial_{h_{\rm x}}+{\rm sign}(l\ell)\partial_{h_{\rm y}}\bigr)^{|l|}\bigl({\rm i}\partial_{h_{\rm x}}+{\rm sign}(l'\ell)\partial_{h_{\rm y}}\bigr)^{|l'|}\,{\rm e}^{-\frac{(w_0\vec{k}_\perp+\vec{h}_\perp)^2}{8}}\nonumber\\
 &\quad \times  F_{N+N'+|l|+|l'|+2(j+j'+1)}\bigl(\bigl|h_{\rm z}-{\rm z}_R(k\hat\kappa)+\ell\tfrac{(w_0\vec{k}_\perp+\vec{h}_\perp)^2}{8}\bigr|,0\bigr)\Big|_{\vec{h}=0}
\end{align}
and
\begin{align}
  ({\cal J}_\ell^-)_{{\cal N},j;{\cal N}',j'}^{{\rm LG}}(k)&={\rm e}^{{\rm i} \ell(\varphi_{l,p}-\varphi_{l',p'})} \bigl[1+{\rm sign}\bigl(\ell(N-N')\bigr)\partial_{h_{\rm z}}\bigr]^{|N-N'|}\, \partial_c^{j+j'} \nonumber\\ 
 &\quad\times 
 \bigl({\rm i}\partial_{h_{\rm x}}+{\rm sign}(l\ell)\partial_{h_{\rm y}}\bigr)^{|l|}\bigl({\rm i}\partial_{h_{\rm x}}-{\rm sign}(l'\ell)\partial_{h_{\rm y}}\bigr)^{|l'|} \nonumber\\  
 &\quad\times \frac{1}{c}F_{|N-N'|+|l|+|l'|}\bigl(|h_{\rm z}-{\rm z}_R(k\hat\kappa)|,\tfrac{(w_0\vec{k}_\perp+\vec{h}_\perp)^2}{8c}\bigr) \Big|_{c=1,\,\vec{h}=0}\,.
\end{align}

\section{Conclusions and Outlook}

In this proceedings, we have determined analytic expressions for the one-loop photon polarization tensor in circularly polarized Hermite- and Laguerre-Gaussian beams, thereby complementing our recent study~\cite{Karbstein:2017jgh} for linearly polarized beams.
Our results are based on a LCFA of the Heisenberg-Euler effective Lagrangian for QED in constant fields, and thus are manifestly limited to slowly varying electromagnetic fields.
More specifically, our results neglect contributions of the following type, 
\begin{equation}
\sim\upsilon^2\Bigl[{\cal O}\bigl(\tfrac{2}{w_0\Omega}\bigr)+{\cal O}\bigl(\tfrac{1}{\tau\Omega}\bigr)+{\cal O}\bigl((\tfrac{\upsilon}{m_e})^2\bigr)+{\cal O}\bigl((\tfrac{\alpha}{\pi})^2\bigr)\Bigr]\,,
\end{equation}
where $\{\omega,|\vec{k}|,\omega',|\vec{k}'|,\Omega\}\lesssim\upsilon\ll m_e$; see Ref.~\cite{Karbstein:2017jgh} for the details.
The first term in the squared brackets arises from the restriction to the (leading order) paraxial approximation, the second one is due to the finite pulse duration, the third one refers to contributions beyond the LCFA, and the last one to contributions from higher loops.

As any paraxial beam can be decomposed into either Hermite- or Laguerre-Gaussian modes, our results are expected to be relevant for the study of optical signatures of QED vacuum nonlinearity in realistically modeled, circularly polarized high-intensity lasers pulses.

\ack 

F.K. is grateful to Matt Zepf for stimulating discussions.
The work is supported by Russian Science Foundation grant No. 17-72-20013.
%The research of E.A.M. is funded by the Tomsk Polytechnic University Competitiveness Enhancement Program grant, CEP-PTI-72/2017.

\end{document}